Lipid Case Studies
Three patients are seen in clinic:
-Patient one is a 40 year old man with hypertension, who also smokes, but has not been previously diagnosed with CHD. His father developed CHD at age 53 years. He is fasting, and the results of his lipids include a total cholesterol concentration of 210 mg/dL, triglycerides of 150 mg/dL, and an HDL cholesterol level of 45 mg/dL. He has a fasting glucose level of 98 mg/dL.

-Patient two is a 60-year old woman with no family history of CHD, who is normotensive and does not smoke, with a total cholesterol concentration of 220 mg/dL, triglycerides of 85 mg/dL, and an HDL cholesterol value of 80 mg/dL. Her fasting glucose level is 85 mg/dL.

-Patient three is a 49 year old man with no personal or family history of CHD, and who is not hypertensive and does not smoke. His fasting total cholesterol level is 260 mg/dL, his triglycerides are 505 mg/dL, his HDL cholesterol is 25 mg/dL, and his glucose level is 134 mg/dL.

QUESTIONS

What is the LDL cholesterol level, as calculated using the Friedewald calculation?
Which patient, if any, should have their LDL cholesterol measured, rather than calculated? Why?
How many known CHD risk factors does each patient have?
Based on what is known, are these patient recommended for lipid therapy (diet or drug) and, if so, on what basis?
Answer #1

Before going in details first we have to learn about normal values of following:

* Total cholesterol = 150-220 mg/dl

* Triglycerides = 60-165 mg/dl---male ; 40-140 mg/dl---female.

* HDL= 35-55 mg/dl---male; 45-65mg/dl---female

* LDL=60-140 mg/dl

* Fasting blood glucose= 60-100mg/dl

* Random blood sugar= 100-140 mg/dl

1) LDL levels are calculated using friedwald calculation method which is as following:

Tg÷5 =VLDL

Total cholesterol - (VLDL+HDL)=LDL

Putting values in above cases #Case one have

Tg = 150mg/dl

HDL= 45mg/dl

TOTAL cholesterol= 210mg/dl

Fasting glucose= 98mg/dl

First we have to calculate VLDL = TG÷5 ;150÷5=30mg/dl

HDL = 45

LDL= TC-(VLDL+HDL); 210-(45+30);210-75=135mg/dl

# Case two values ;

Total cholesterol=220 mg/dl

Triglycerides=85mg/dl

HDL=80mg/dl

Fasting Blood glucose=85mg/dl

VLDL= Tg÷5;85÷5=17mg/dl

LDL= TC-(VLDL-HDL): 220-97=123mg/dl

#case three:

VLDL= 505÷5=101mg/dl

LDL= 260-126=34mg/dl

2) Patient three with high TG levels are advised to go for LDL measurements by laboratory because higher tg levels decreases the reliablity of friedwald calculation.

3)Coronary heart diseases risk factors are:

*High LDL conc.

*Smoking

*Family history

* High BP

*Postmenopausal women

* Men above age of 45

* Low HDL

Risk factors in #case one: Hypertension,smoking,positive family history.

#case two:

Age ,high concentration of tg

# case three:

Age above 45,high tg concentration,

4)Lipid therapy a patient with concentration of more than 100mg/dl LDL is indicated for lipid therapy. So case one and two are advised for lipid therapy and case three is advised for further LDL measurements which are more specific and reliable.